# Harnessing The Collective Wisdom: Fusion Learning Using Decision Sequences From Diverse Sources

### v2

@@ 2 IRT: integrative FDR control using binary decision sequences
In the following sections we present IRT, an Integrative Ranking and Thresholding framework that involves three steps. In Step 1, IRT utilizes the binary decision sequence $\bm{\delta}_{j}$ from study $j$ to construct a measure of evidence against the null hypotheses. In Step 2, this evidence is aggregated into a discriminatory measure such that for each null hypothesis $H_{0i}$, a large aggregated evidence implies stronger evidence against $H_{0i}$. In the final Step 3, the rankings of aggregated evidence are used to determine a cutoff for FDR control. In what follows, we describe each of these steps in detail while Algorithm 1 summarizes the discussion below.
2.1 Step 1: Evidence Construction
IRT uses the information in $\mathcal{D}_{j}$ to construct an evidence index $e_{ij}$ as follows:
$e_{ij}=w_{j}\dfrac{\delta_{ij}}{\|\bm{\delta}_{j}\|_{0}\vee 1}{\text{~{}where~{}}w_{j}=\frac{m_{j}}{\alpha_{j}}},\quad\forall i\in\mathcal{M}_{j}.$ (1)
The evidence weights $w_{j}$ in Equation (1) capture the relative importance of a rejection across the $d$ studies and, hence, play a key role in differentiating among them. For study $j$, the evidence weight is distributed evenly across the rejected hypotheses. Moreover, this weight is higher if study $j$ has tested more hypotheses (larger $m_{j}$) and more conservative (smaller $\alpha_{j}$). This is in perfect accordance with our intuition since the evidence needed to reject a hypothesis increases as the number of hypotheses tested increases or the target FDR level decreases. For example, if only one hypothesis is tested then rejecting that null hypothesis when the $p-$value is below $\alpha$ guarantees FDR is controlled at level $\alpha$. However, if more than one hypothesis is tested then simply rejecting null hypotheses with $p-$value at most $\alpha$ does not guarantee FDR control at level $\alpha$.
The evidence indices $\mathbf{e}_{j}=\{e_{ij}\}_{i\in\mathcal{M}_{j}}$ in Equation (1) are related to $e-$values for hypothesis testing, which serve as a natural counterpart to the widely adopted $p-$values. A non-negative random variable $e$ is an “$e-$value” [^2] We will use the notation ‘$e$’ to denote both the random variable and its realized value. if $\mathbb{E}[e]\leq 1$ under the null hypothesis. A large $e-$value provides evidence against the null hypothesis. See Vovk and Wang (2021) for a background on $e-$values for hypothesis testing. Recently, Wang and Ramdas (2022); Ren and Barber (2024) introduce the concept of generalized $e-$values which are defined as follows:
Definition 1 (Generalized $e-$values).
@@ 2.1 Step 1: Evidence Construction
Let $\mathbf{e}=\{e_{1},\dots,e_{m}\}$ be a collection of random variables associated with the null hypotheses $H_{01},\dots,H_{0m}$. We say $\mathbf{e}$ is a set of generalized $e-$values if $\sum_{i\in\mathcal{H}_{0}}\mathbb{E}[e_{i}]\leq m$.
Theorem 1 establishes that the evidence indices in Equation (1) are generalized $e-$values.
@@
Suppose study $j$ controls FDR at level $\alpha_{j}$. Then, $\mathbf{e}_{j}$ is a set of generalized $e-$values associated with $\mathcal{H}_{0j}=\{H_{0i}:i\in\mathcal{M}_{j}\}$.
Proof of Theorem 1..
Based on the evidence construction in (1), we have
$\sum_{i\in\mathcal{H}_{0j}}\mathbb{E}[e_{ij}]$ $=\dfrac{m_{j}}{\alpha_{j}}\cdot\mathbb{E}\left[\dfrac{\sum_{i\in\mathcal{H}_{0j}}\delta_{ij}}{\|\bm{\delta}_{j}\|_{0}\vee 1}\right]=\dfrac{m_{j}}{\alpha_{j}}\cdot\text{FDR}(\bm{\delta}_{j})\leq m_{j},$
To provide some intuitions behind the construction of evidence index in Equation (1), we discuss Step 1 under the light of an important special case of testing a single hypothesis. Suppose $m=d=1$ and we are testing $H_{01}$. Given a valid level $\alpha_{1}$ testing procedure $\delta$, the corresponding evidence index:
$e_{11}=(1/\alpha_{1})\cdot\mathbb{I}(\delta\ \text{rejects}\ H_{01})$ (2)
is the most powerful $e-$value that is equivalent to this testing procedure at threshold $1/\alpha_{1}$. The evidence index in Equation (2) is also known as the all-or-nothing bet (Shafer, 2021) against $H_{01}$ and provides another intuitive interpretation of the evidence indices in Equation (1) as follows: if the $j$-th study controls FDR at level $\alpha_{j}$ then $\mathbf{e}_{j}$ are scaled all-or-nothing bets against the null hypotheses in $\mathcal{M}_{j}$ where the scaling $\frac{m_{j}}{\|\bm{\delta}_{j}\|_{0}\vee 1}$ ensures that if all the null hypotheses corresponding to the non-zero bets $\{i\in\mathcal{M}_{j}:e_{ij}\neq 0\}$ are true then the probability that the sum of those non-zero bets exceed $m_{j}/\alpha_{j}$ is at most $\alpha_{j}$. In Section 4.1 we provide another motivation for the evidence indices and show that $\mathbf{e}_{j}$ naturally arise as building blocks of several popular aggregation and derandomization procedures, such as the derandomized knockoff of Ren and Barber (2024) and the $e-$value aggregation procedure of Li and Zhang (2023).
2.2 Step 2: Evidence Aggregation
IRT aggregates the evidence indices $\mathbf{e}_{j}$ across the studies as follows:
${e}_{i}^{\texttt{agg}}=\dfrac{m}{\sum_{j=1}^{d}m_{j}}\sum_{j=1}^{d}e_{ij}\cdot\mathbb{I}(i\in\mathcal{M}_{j}),\quad\forall i\in\mathcal{M}.$ (3)
In Equation (3), $e_{i}^{\tt agg}$ represents the aggregated evidence across all studies that test hypothesis $H_{0i}$. When each study tests all the $m$ hypotheses and $m_{j}=m$, then $e_{i}^{\tt agg}$ is the arithmetic mean of the $d$ evidence indices corresponding to hypothesis $i$, which essentially dominates any symmetric aggregation function by Proposition 3.1 of Vovk and Wang (2021). However, when $m_{j}$ are different, the aggregation scheme in Equation (3) is a natural counterpart to the arithmetic mean of the $d$ evidence indices. Furthermore, in this setting Theorem 2 establishes that $\mathbf{e}^{\texttt{agg}}=\{e_{i}^{\texttt{agg}}\}_{i\in\mathcal{M}}$ is a set of generalized $e-$values associated with $\mathcal{H}_{0}$.
@@ 2.2 Step 2: Evidence Aggregation
Suppose that each study $j$ controls FDR at level $\alpha_{j}$. Then, $\mathbf{e}^{\texttt{agg}}$ is a set of generalized $e-$values associated with $\mathcal{H}_{0}$.
Proof of Theorem 2..
We have
$\sum_{i\in\mathcal{H}_{0}}\mathbb{E}[e_{i}^{\texttt{agg}}]$ $=\dfrac{m}{\sum_{j=1}^{d}m_{j}}\sum_{i\in\mathcal{H}_{0}}\sum_{j=1}^{d}\dfrac{m_{j}}{\alpha_{j}}\cdot\mathbb{E}\left[\dfrac{\delta_{ij}}{\|\bm{\delta}_{j}\|_{0}\vee 1}\right]\cdot\mathbb{I}(i\in\mathcal{M}_{j})\leq\dfrac{m}{\sum_{j=1}^{d}m_{j}}\sum_{j=1}^{d}\dfrac{m_{j}}{\alpha_{j}}\cdot{\rm FDR}(\bm{\delta}_{j})\leq m,$
which completes the proof. ∎
In Section 3.2 we discuss an alternative evidence aggregation scheme under some additional assumptions on the data generating process for each study.
2.3 Step 3: FDR control
In the context of multiple testing with $e-$values, Wang and Ramdas (2022) have proposed the $e-$BH procedure that controls the FDR at level $\alpha$ even under unknown arbitrary dependence between the generalized $e-$values. The $e-$BH procedure is related to the well-known Benjamini-Hochberg (BH) procedure (Benjamini and Hochberg, 1995) and can be summarized as follows: let $\mathbf{e}=\{e_{i}\}_{i\in\mathcal{M}}$ be a set of generalized $e-$value associated with the null hypotheses $\mathcal{H}_{0}$. Denote $e_{(1)}\geq\ldots\geq e_{(m)}$ as the ordered $e-$values from largest to smallest. The rejection rule of $e-$BH is given by $\delta_{i}=\mathbb{I}\left(e_{i}\geq{m}/{\alpha k_{\alpha}}\right)$ for all $i\in\mathcal{M}$, where the threshold is chosen as
$k_{\alpha}=\max\left\{i\in\mathcal{M}:e_{(i)}\geq\dfrac{m}{i\alpha}\right\}.$ (4)
In step 3, IRT uses $\mathbf{e}^{\tt agg}$ as an input for the $e-$BH procedure to get the final rejection set and, in conjunction with Theorem 2 and Theorem 2 of Wang and Ramdas (2022), the IRT framework controls FDR at level $\alpha$ under unknown arbitrary dependence between the aggregated evidence indices $\mathbf{e}^{\tt agg}$. We summarize the aforementioned three steps in Algorithm 1.
Algorithm 1 Integrative Ranking and Thresholding (IRT)
Inputs: $\mathcal{D}_{j}=\{\bm{\delta}_{j},\alpha_{j},\mathcal{M}_{j}\}$ for $j=1,\ldots,d$ studies, designated FDR level $\alpha\in(0,1)$.
1:// (Step 1) Evidence construction
2:Construct $e_{ij}={m_{j}\delta_{ij}}/\{\alpha_{j}(\|\bm{\delta}_{j}\|_{0}\vee 1)\}$ for each $(j,i)\in[d]\times\mathcal{M}_{j}$.
3:// (Step 2) Evidence aggregation
4:Calculate $e_{i}^{\tt agg}$ as in Equation (3) or ${e}_{i}^{\tt agg*}$ as in Equation (7) if conditions of Theorem 3 hold for each $i\in\mathcal{M}$.
5:// (Step 3) FDR control with $e-$BH procedure
6:if conditions of Theorem 3 hold then
7:     Choose the threshold as $k_{\alpha}=\max\left\{i\in\mathcal{M}:e^{\tt agg*}_{(i)}\geq\dfrac{m}{i\alpha}\right\}$and reject all $H_{0i}$ with $e^{\tt agg*}_{i}\geq m/(\alpha k_{\alpha})$.
8:else
9:     Choose the threshold as $k_{\alpha}=\max\left\{i\in\mathcal{M}:e^{\tt agg}_{(i)}\geq\dfrac{m}{i\alpha}\right\}$and reject all $H_{0i}$ with $e^{\tt agg}_{i}\geq m/(\alpha k_{\alpha})$.
10:end if
3 IRT: more powerful evidence aggregation via multiplication

### v3

@@ 2.2 The IRT framework
The proposed IRT framework involves three steps. In Step 1, IRT utilizes the binary decision sequence $\bm{\delta}_{j}$ from study $j$ to construct a measure of evidence against the null hypotheses. In Step 2, this evidence is aggregated into a discriminatory measure such that for each null hypothesis $H_{0i}$, a large aggregated evidence implies stronger evidence against $H_{0i}$. In Step 3, the aggregated evidence scores are used to produce a final set of discoveries with guaranteed FDR control. In what follows, we describe each of these steps in detail.
Step 1: Evidence Construction - To build intuition, we first consider the familiar setting where study $j$ reports decisions $\bm{\delta}_{j}$ from applying the Benjamini-Hochberg (BH) procedure (Benjamini and Hochberg, 1995) to its corresponding $p$-values, $\bm{p}^{*}_{j}=(p^{*}_{ij}:i\in\mathcal{M}_{j})$, at FDR level $\alpha_{j}$. The information available to IRT is the triplet $\mathcal{D}_{j}=\{\bm{\delta}_{j},\alpha_{j},\mathcal{M}_{j}\}$, which notably excludes $\bm{p}^{*}_{j}$.
Define $t_{j}=(\alpha_{j}/m_{j})\|\bm{\delta}_{j}\|_{0}$, which is fully determined by the available information in $\mathcal{D}_{j}$. While the true $p$-value $p^{*}_{ij}$ is unobserved, based on the mechanics of the BH procedure we know that $p^{*}_{ij}\leq t_{j}$ for any rejected hypothesis ($\delta_{ij}=1$), and $p^{*}_{ij}>t_{j}$ for any non-rejected one ($\delta_{ij}=0$). This allows us to define a conservative $p$-value:
$p_{ij}=t_{j}\cdot\mathbb{I}(\delta_{ij}=1)+1\cdot\mathbb{I}(\delta_{ij}=0).$ (1)
By construction, $p^{*}_{ij}\leq p_{ij}$, meaning $p_{ij}$ is a valid, but conservative, $p$-value for $H_{0i}$. While these conservative $p$-values could be used in traditional meta-analysis, an alternative and increasingly popular approach is to transform evidence into e-values (Vovk and Wang, 2021). An $e$-value is a non-negative random variable ee [^2] We will use the notation ‘$e$’ to denote both the random variable and its realized value. with the property that $\mathbb{E}[e]\leq 1$ under the null hypothesis; a large $e$ indicates strong evidence against the null. A key advantage of $e$-values over $p$-values is their ease of aggregation. In fact, Vovk et al. (2022) show that admissible methods for combining $p$-values under arbitrary dependence essentially operate by first converting $p$-values into $e$-values and then averaging them. In the context of our BH example, if the original $p$-values $\{p^{*}_{ij}\}$ are independent, it can be shown that transforming our conservative $p$-values via
$e_{ij}=\begin{cases}\pi_{j}p_{ij}^{-1}&\text{if }\delta_{ij}=1\\ 0&\text{if }\delta_{ij}=0\end{cases},$ (2)
where $\pi_{j}=|\mathcal{H}_{0j}|/m_{j}$, yields a valid $e$-value for each hypothesis $H_{0i}$ (see Lemma 1 in Supplement B).
The central insight of our work is that this principle extends far beyond the BH procedure and does not require the FDR procedure to use $p$-values at all. To generalize this idea to decisions from any FDR-controlling method (such as knockoffs (Barber and Candès, 2015) or covariate-powered methods (Ignatiadis and Huber, 2021)) using arbitrarily dependent test statistics, we shift our perspective from individual $e$-values to compound $e$-values (Wang and Ramdas, 2022; Ren and Barber, 2024; Ignatiadis et al., 2024).
Definition 1 (Compound $e$-values).
@@
Let $\bm{e}=\{e_{1},\dots,e_{m}\}$ be a collection of random variables associated with the hypotheses $H_{01},\dots,H_{0m}$. We say $\bm{e}$ is a set of compound $e$-values if $\sum_{i\in\mathcal{H}_{0}}\mathbb{E}[e_{i}]\leq|\mathcal{H}_{0}|$.
With this concept, we now define the general evidence construction for IRT, which operates directly on the triplet $\mathcal{D}_{j}$:
$e_{ij}=w_{j}\frac{\delta_{ij}}{\max(\|\bm{\delta}_{j}\|_{0},1)},\text{ where the evidence weight }w_{j}=\frac{m_{j}}{\alpha_{j}},~\forall i\in\mathcal{M}_{j}.$ (3)
The term $\delta_{ij}/\max(\|\bm{\delta}_{j}\|_{0},1)$ distributes the “evidence weight” of study $j$ evenly across its rejected hypotheses. The weight $w_{j}$ assigns greater importance to rejections from studies that are larger (more hypotheses $m_{j}$) or more conservative (smaller FDR level $\alpha_{j}$). This general construction is the cornerstone of our framework, leading to our main theoretical result for this step.
@@
Suppose study $j$ controls FDR at level $\alpha_{j}$. Then the collection of evidence indices $\bm{e}_{j}=\{e_{ij}\}_{i\in\mathcal{M}_{j}}$ from Equation (3) is a set of compound $e$-values associated with the null hypotheses in $\mathcal{H}_{0j}$.
In Section D.1 of the supplement, we show that these evidence indices also naturally arise as building blocks of several popular aggregation and derandomization procedures, such as those of Ren and Barber (2024) and Li and Zhang (2023). Section E explores how compound $e$-values can be derived from decisions that control alternative notions of Type I error.
Step 2: Evidence Aggregation - Given the compound $e$-values from each study, IRT aggregates the evidence indices $\mathbf{e}_{j}$ across the studies as follows:
${e}_{i}^{\texttt{agg}}=\dfrac{1}{d}\sum_{j=1}^{d}\Bigl{\{}e_{ij}~\mathbb{I}(i\in\mathcal{M}_{j})+\mathbb{I}(i\notin\mathcal{M}_{j})\Bigr{\}},\quad\forall i\in\mathcal{M}.$ (4)
When each study tests all the $m$ hypotheses and $m_{j}=m$, then $e_{i}^{\tt agg}$ is the arithmetic mean of the $d$ evidence indices corresponding to hypothesis $i$. However, when $m_{j}$ are different, the aggregation scheme in Equation (4) sets $e_{ij}=1$ whenever $i\notin\mathcal{M}_{j}$, which is a valid $e$-value [^3] Note that in our framework, we treat the $e$-values for hypotheses in $\mathcal{M}\setminus\mathcal{M}_{j}$ as missing completely at random (Rubin, 1976). Here “ $\setminus$ ” is the usual set difference operator.. This aggregation preserves the compound $e$-value property as the next Theorem shows.
@@
Suppose that each study $j$ controls FDR at level $\alpha_{j}$. Then, $\bm{e}^{\texttt{agg}}=\{e_{i}^{\texttt{agg}}:i\in\mathcal{M}\}$ is a set of compound $e$-values associated with $\mathcal{H}_{0}$.
The intuition for this choice is that for standard $e$-values, simple averaging is known to dominate any other symmetric aggregation function under arbitrary dependence (Vovk and Wang, 2021; Wang, 2024). While we are working with compound $e$-values of a specific form, this provides strong heuristic justification for our approach. Whether simple averaging is formally admissible in this specific setting remains an interesting open question.
Step 3: FDR control - Once the aggregated evidence indices $\bm{e}^{\,\texttt{agg}}$ are constructed, there are two primary paths to obtain a final set of rejections, depending on the available data. The most direct approach is to apply the e-BH procedure (Wang and Ramdas, 2022) to the set of aggregated compound $e$-values, $\bm{e}^{\,\texttt{agg}}$. Specifically, denote $e_{(1)}\geq\ldots\geq e_{(m)}$ as the ordered $e$-values from largest to smallest. The rejection rule of e-BH is given by $\delta_{i}=\mathbb{I}\{e_{i}\geq{m}/(\alpha k_{\alpha})\}$ for all $i\in\mathcal{M}$, where the threshold is chosen as $k_{\alpha}=\max\{i\in\mathcal{M}:e_{(i)}\geq{m}/({i\alpha})\}$ with the convention that $\max{(\emptyset)}=0$. This method guarantees FDR control at the desired level $\alpha$ under arbitrary dependence structures among the $e$-values. However, a key limitation of this approach is that it cannot make rejections if the target FDR level $\alpha$ is more stringent than that of any individual study, i.e., if $\alpha<\min_{j\in[d]}\alpha_{j}$. This may seem counterintuitive for meta-analysis, but it is not a flaw of our method. Rather, it is an intrinsic property of any procedure that aggregates evidence derived solely from binary decisions. Without additional information or assumptions, one cannot generate evidence stronger than the strongest input. A more detailed discussion of this property is provided in Section A of the supplement.
Alternatively, IRT provides a powerful building block for integrative analyses when heterogeneous data types are available. Consider a common meta-analysis setting where, in addition to the $d$ studies providing binary decisions, we have access to a separate, independent study that reports a full set of $p$-values, $\bm{p}=\{p_{1},\ldots,p_{m}\}$. We can fuse this information with our aggregated compound $e$-values, $\bm{e}^{\texttt{agg}}$, using the ep-BH procedure (Ignatiadis et al., 2024). This method treats the $e$-values as unnormalized, data-driven weights for the $p$-values. Specifically, one first computes a set of re-weighted $p$-values, $p^{\prime}_{i}=\min(p_{i}/e_{i}^{\texttt{agg}},1)$, and then applies the standard BH procedure to this new set $\{p^{\prime}_{i}:i\in\mathcal{M}\}$. The primary benefit of this hybrid approach is its potential for greater statistical power compared to analyzing each data source separately. That is, the ep-BH procedure can yield more discoveries than either: (1) applying the standard BH procedure to the independent $p$-values $\bm{p}$ alone, or (2) applying the e-BH procedure to the aggregated $e$-values $\bm{e}^{\texttt{agg}}$ alone. Section 3 illustrates this with numerical examples.
3 Numerical illustrations
We illustrate the utility of IRT for meta-analysis under the setting where $d_{1}$ studies report $p$-values while $d_{2}$ studies report their binary decision sequences. Tang et al. (2014) discuss $p$-value imputation techniques for meta-analysis under this setting but assume that the $d=d_{1}+d_{2}$ studies are independent. In the next two examples, we demonstrate that when the studies are dependent, in a sense that is discussed subsequently, IRT provides a powerful strategy for pooling inferences in this scenario.
Suppose, without loss of generality, the first $d_{1}$ studies report their raw $p$-values for each of the $m$ hypotheses while the remaining $d_{2}=d-d_{1}$ studies report the triplet $\mathcal{D}_{j},j=d_{1}+1,\ldots,d$, where the corresponding decision sequences $\bm{\delta}_{j}$ are obtained from the BH procedure with FDR control level $\alpha_{j}=0.01$. We set $m=1000$ and consider testing $H_{0i}:\mu_{i}=0~vs~H_{1i}:\mu_{i}\neq 0$, where $\mu_{i}\stackrel{{\scriptstyle\rm i.i.d.}}{{\sim}}0.95~\delta_{\{0\}}+0.025~\mathcal{N}(3,1)+0.025~\mathcal{N}(-3,1)$ and $\delta_{\{a\}}$ denotes a point mass at $a$. For each hypothesis $i$, the test statistics $\bm{X}_{i}=(X_{ij}:j\in[d])\stackrel{{\scriptstyle ind.}}{{\sim}}\mathcal{N}_{d}(\mu_{i}\bm{1}_{d},\bm{\Sigma})$ where $\bm{\Sigma}=\begin{pmatrix}\bm{\Sigma}_{d_{1}}&\bm{0}_{d_{1}\times d_{2}}\\ \bm{0}_{d_{2}\times d_{1}}&\bm{\Sigma}_{d_{2}}\end{pmatrix}$ so that the $d_{1}$ and $d_{2}$ studies are independent of each other. We set $\bm{\Sigma}_{d_{k}}=\rho_{k}\bm{1}_{d_{k}}\bm{1}_{d_{k}}^{T}+(1-\rho_{k})\bm{I}_{d_{k}}$ for $k\in\{1,2\}$ and $\rho\in(0,1)$. The raw $p$-values are computed using the standard two-sided $Z$-test formula, $p_{ij}=2\Phi(-|X_{ij}|)$, where $\Phi$ is the distribution function for standard normal. The following four methods are evaluated for integrative inference at FDR level $\alpha$: (i) Cauchy $d_{1}$, a baseline which derives the pooled $p$-values from the first $d_{1}$ studies using the Cauchy combination test statistic (Liu and Xie, 2020) followed by a BH correction; (ii) Cauchy, an idealized benchmark which is similar to Cauchy $d_{1}$ but derives the pooled $p$-values from all $d$ studies, (iii) Cauchy + Imputed, that first imputes the $p$-values for the $d_{2}$ studies using Equation (1), then pools all $p$-values using the Cauchy combination test statistic and finally applies the BH correction, (iv) Cauchy + IRT, which is the hybrid approach. We first use the Cauchy combination test to produce a single pooled $p$-value vector, $\bm{p}$, from the $d_{1}$ studies. We then apply IRT to the $d_{2}$ studies to generate aggregated $e$-values, $\bm{e}^{\texttt{agg}}$. Finally, we apply the ep-BH procedure (Ignatiadis et al., 2024) to the pairs $\{\bm{p},\bm{e}^{\texttt{agg}}\}$. This procedure controls FDR because the block diagonal structure of $\bm{\Sigma}$ ensures $\bm{p}$ and $\bm{e}^{\texttt{agg}}$ are independent. The performances of these four methods are compared with the quality of inference obtained from a single study that applies the BH procedure on the $p$-values from the first study for FDR control at level $\alpha_{1}$.
Example 1.
We fix $d_{1}=2,~d_{2}=3,~\rho_{k}=\rho$ and vary $\rho\in\{0,0.1,0.3,0.5,0.7,0.9\}$. Across $2000$ Monte-Carlo (MC) repetitions of this data generating scheme, Figure 1 reports the average FDP and ETP of the five different methods for integrative inference at the FDR level $0.5\%$. All methods for integrative inference control FDR at $0.5\%$ while the single study controls it at its designated FDR level $1\%$. Unsurprisingly, Cauchy, overall, has the highest power across almost all values of $\rho$ as it uses the $p$-value information from all $d$ studies for integrative inference. In contrast, inferences that rely on a conservative $p$-value imputation method have the least power. This is followed by Cauchy $d_{1}$ which fuses inference across the first $d_{1}$ studies. When $\rho$ is high, pooled inferences from Cauchy $d_{1}$ are less powerful than those from a single study. This is expected since Cauchy $d_{1}$ controls FDR at a more stringent level and pooling inferences across highly dependent studies may not lead to substantially more true positives than what can be learned from a single study. The key comparison is between Cauchy $d_{1}$ and Cauchy + IRT. Across all levels of correlation, Cauchy + IRT is uniformly more powerful. This result directly demonstrates the value of the information contained in the binary decisions. By transforming them into compound $e$-values, IRT allows us to extract meaningful evidence and achieve greater statistical power than an analysis that discards this information.
Figure 1: FDP and ETP comparison for Example 1.
Example 2.
We now examine a setting with asymmetric dependence, where the $p$-value studies are correlated (with correlation $\rho_{1}$) while the binary-decision studies are independent ($\rho_{2}=0$). We continue to take $d_{1}=2,~d_{2}=3$ but set $\alpha=\alpha_{j}=0.01$ for $j\in[d]$. Figure 2 reports the average FDP and ETP for various methods across 2000 MC repetitions. While all methods control FDR at $1\%$, we find that across all values of $\rho_{1}$, Cauchy + IRT is more powerful than Cauchy + Imputation, Cauchy $d_{1}$ and the Single Study.
Figure 2: FDP and ETP comparison for Example 2.
Additional numerical experiments and a real data analysis illustrating the empirical performance of IRT are, respectively, presented in sections F and G of the Supplement.
4 Concluding remarks
IRT is a general framework for fusion learning in multiple testing that operates on the binary decision sequences available from diverse studies and conducts integrative inference on the common parameter of interest. For meta-analysis involving dependent studies, IRT provides a powerful alternative for fusing inferences when some studies report $p$-values while the rest reveal only the rejections under a pre-specified FDR level. Section B of the supplement proposes IRT and IRT H which are powerful alternatives to IRT under additional assumptions on the data generating process for each study.
[…]
Supplementary material
This supplement is organized as follows: Section A discusses the impact of target FDR levels on the power of IRT. In Section B we present IRT that relies on an alternative scheme for evidence aggregation. The proofs of all theoretical results in the paper are presented in Section C. In Section D we present additional insights about the IRT framework. In particular, we show that (i) the evidence indices in Equation (3) are connected to some existing aggregation and derandomization procedures (Section D.1), and (ii) prove that IRT guarantees asymptotic FDR control if some studies control their FDR asymptotically (Section D.2). Section E extends IRT to alternative forms of Type I error control. Additional numerical studies and a real data application are presented in sections F and G, respectively.
Appendix A Impact of $\alpha$ on the power of IRT